OpenStudy (anonymous):

Lucas recorded his lunch expenditure each day for one week in the table below. Day Sunday Monday Tuesday Wednesday Thursday Friday Saturday Expenditure $4.85 $5.10 $5.50 $4.75 $4.50 $5.00 $6.00 Find the standard deviation, and variance of Lucas’ lunch expenditures. Round to the nearest thousandth

OpenStudy (anonymous):

@campbell_st @CaityxD @genius12 please help me?!

OpenStudy (campbell_st):

well if you don't have a calculator with a stats mode you need to 1. find the mean \[\mu = \frac{4.85 + 5.10+5.50+4.75+4.50+5.00+6.00}{7}\] the use the standard deviation formula \[\sigma = \sqrt{\frac{1}{N}[\sum_{i = 1}^{N}(x_{i} - \mu)]}\] \[\sigma = \sqrt{\frac{1}{7}[(4.85 - \mu)^2 + (5.10 - \mu)^2 + (5.1 - \mu)^2 + (4.75 - \mu)^2 + (4.5-\mu)^2 +(5 - \mu)^2+(6 - \mu)^2}\] or you can google an online standard deviation calculator an alternative is